Let X be a random variable with the following probability distribution:

 

X

-3

6

9

P(X=x)

1/6

1/2

1/3

 

If the function of random variable is defined as f(X) = (2X+1)2, find µg(X).

 


1
Expert's answer
2021-08-06T11:37:12-0400

"E[(2X+1)^2]=(-3*2+1)^2*\frac{1}{6}+(6*2+1)^2*\frac{1}{2}+(9*2+1)^2*\frac{1}{3}="

"=\frac{1254}{6}=209."